OpenStudy (anonymous):

in and equation like, 6x-7=4x+23 Since one side is subtract and one is addition do i subtract from seven from both side or would i add?

OpenStudy (anonymous):

? is that some kind of sick joke because that didnt help me much

OpenStudy (anonymous):

hello

OpenStudy (ash2326):

@smarie96 We have \[6x-7=4x+23\] If we want x on the left side, we have to add 7 both sides and subtract 4x from both sides. \[6x-7+7=4x+23+7\] \[6x+0=4x+30\] Subtract 4x from both sides \[6x-4x=4x+30-4x\] \[2x=30\] Do you understand this?

OpenStudy (ash2326):

@smarie96 Are you here?

OpenStudy (anonymous):

yes thank you

OpenStudy (ash2326):

There is one more way we could bring x on the right side. In that case we'll subtract 26 from both sides

OpenStudy (anonymous):

okay so if both sides are addition i just subtract both sides?

OpenStudy (ash2326):

yes

OpenStudy (anonymous):

so for example if its like x+7=2x+14 i would subtract seven from both sides so it would look like x+7-7=2x+14-7 x+0=2x+7 X+0-2=7 x=7?

OpenStudy (ash2326):

First step is right \[x=2x+7\] Now subtract 2x from both the sides

OpenStudy (anonymous):

so the answer is 7 right? so all i need to change is subtract and not add the 2x

OpenStudy (ash2326):

\[x-2x=2x-2x+7\] \[-x=7\] \[x=??\]